OpenStudy (anonymous):

use point slope form to write the equation of the line that has a slope of 4 and papssed though the point (-5,-3)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(-5\quad ,&-3) \end{array} \\\quad \\ slope = m=4 \\ \quad \\ y-y_1=m(x-x_1)\quad \textit{plug in your values and solve for "y"}\)